OpenStudy (anonymous):

find the equation of the parabola that passes through the points (2,0),(3,-2),(1,-2) I know the equation you have to write is y=ax^2+bx+c you take x and y and plug them into that equation. Therefore, it looks like this Set 1: 4a+2b+c=0 Set 2: 9a+3b+c=-2 Set 3: a+b+c=-2 now i have to know how to find a,b, and c to make that equation y=ax^2+bx+c but i'm not sure how

OpenStudy (superdavesuper):

3 eqns n 3 unknown - a,b,c so it can definitely be solved. u can start by subtract Set 2 from Set 1 n end up with a new eqn w only a and b...

OpenStudy (superdavesuper):

Hint: Set 1: 4a+2b+c=0 - Set 2: 9a+3b+c=-2 --------------------- -5a - b = 2 so b=-5a-2 put that into Set 3...